# On the non-hermitian Kitaev chain

### v1

@@ VII Discussion
We studied the non-hermitian Kitaev chain for general complex parameters, pushing analytical methods as far as possible. We now discuss our results, by means of an example. We use the chain with parameters $m=2/5$, $t_{1}=2+i$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$ for this purpose. One of the main results we obtained in this paper, is the characterisation of the eigenvalues of the infinite size system, in terms of three Vieta Eqs. (31), for $\lambda$ which we repeat here for convenience,
@@
where $\kappa$ and $s$ are in general complex. The eigenvalue curves $\lambda$ are obtained by varying $0\leq\alpha<2\pi$, but only those $\lambda$ that correspond to solutions that satisfy $1/|s|\leq|\kappa|^{2}\leq|s|$ are actual eigenvalues of the model as explained in Sec. V.
@@
The reason why having a concise characterisation for the infinite system size model is useful, is that (as is well known) obtaining the eigenvalues for large, non-hermitian systems showing skin-effect is numerically unstable. We illustrate this using the same example, by plotting the eigenvalues, as obtained using machine precision diagonalisation for system sizes $L=100$, $L=200$, $L=400$ and $L=800$. The results are shown in Fig. 7 as the green dots, together with the infinite system size results.
@@
Figure 7: Finite size eigenvalues for the model with parameters $m=2/5$, $t_{1}=2+i$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The black lines correspond to actual eigenvalues in the limit $L\rightarrow\infty$. The different panels show the numerically obtained eigenvalues (using machine precision) for the sizes $L=100$, $L=200$, $L=400$ and $L=800$, showing the instability of the algorithm (green dots).
@@
Despite the fact that non-hermitian one-dimensional models have been studied in great detail, it would be interesting to apply the method to characterise the eigenvalues in the thermodynamic limit to other, more complicated systems. Here, one can think of models with larger unit cells and models with longer range interactions. Obviously, this will lead to higher order equations and more complicated expressions.
@@ Appendix A Bistritz polynomials for complex parameters
In this appendix, we give the Bistritz polynomials $T_{i}(x)$ (with $4\geq i\geq 1$) in the general case, that is for complex parameters. For details on the Bistritz algorithm, we refer to Bi2002 .
The polynomial $T_{4}(x)$ reads

### v2

@@ VII Analysis of an example
We studied the non-hermitian Kitaev chain for general complex parameters, pushing analytical methods as far as possible. We now summarise our results, by means of an example. We use the chain with parameters $m=2/5$, $t_{1}=2+i$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$ for this purpose. One of the main results we obtained in this paper, is the characterisation of the eigenvalues of the infinite size system, in terms of three Vieta Eqs. (37), for $\lambda$ which we repeat here for convenience,
@@
where $\kappa$ and $s$ are in general complex. The eigenvalue curves $\lambda$ are obtained by varying $0\leq\alpha<2\pi$, and obtaining the solution of the three Vieta Eqs. (37). A priori, the solutions of these equations are continuous and smooth as a function of the parameter $\alpha$. Importantly, only those $\lambda$ that correspond to solutions that satisfy $1/|s|\leq|\kappa|^{2}\leq|s|$ are actual eigenvalues of the model as explained in Sec. V. This explains the observed branched structure of the actual eigenvalues.
@@
We saw that having a concise characterisation of the eigenvalues of the infinite system size explains the structure of the curves the eigenvalues lie on. There is another reason why having a concise characterisation for the infinite system size model is useful, apart from that its interesting in its own right. Namely, as is well known, obtaining the eigenvalues for large, non-hermitian systems showing skin-effect is often numerically unstable. We illustrate this using the same example, by plotting the eigenvalues, as obtained using machine precision diagonalisation for system sizes $L=100$, $L=200$, $L=400$ and $L=800$. The results are shown in Fig. 7 as the green dots, together with the infinite system size results.
@@
Figure 7: Finite size eigenvalues $\lambda$ for the model with parameters $m=2/5$, $t_{1}=2+i$, $t_{2}=1$, $d_{1}=-d_{2}=\sqrt{3}$. The black lines correspond to actual eigenvalues in the limit $L\rightarrow\infty$. The different panels show the numerically obtained eigenvalues (using machine precision) for the sizes $L=100$, $L=200$, $L=400$ and $L=800$, showing the instability of the algorithm (green dots).
@@
VIII Discussion
We studied the non-hermitian Kitaev model for arbitrary, complex parameters. By using a novel method, we obtained a concise characterisation of the eigenvalues in the thermodynamic limit for arbitrary complex parameters. Using this method, we explained the branched structure of the eigenvalues in the complex plane. In addition, we used the solution to obtain for which parameters, the model exhibit a skin effect, and if so, to which eigenvalues this pertains. For real parameters, we obtained this in full, while for arbitrary complex parameters, we obtained sufficient conditions, which we believe are also necessary. We fully characterised the parameters for which the model exhibits a zero mode. Finally, we discussed the stability issues that arise when one tries to numerically obtain the eigenvalues for large systems.
Despite the fact that non-hermitian one-dimensional models have been studied in great detail, it would be interesting to apply our method to characterise the eigenvalues in the thermodynamic limit to other, more complicated systems. One direct extension of the model we consider here, is to add terms that interpolate between open and periodic boundary conditions. Considering different models, one can think of systens with larger unit cells and/or longer range interactions. Obviously, this will lead to higher order equations and more complicated expressions.
@@ Appendix A Bistritz polynomials for complex parameters
In this appendix, we give the Bistritz polynomials $T_{i}(x)$ (with $4\geq i\geq 1$) in the general case, that is for complex parameters. These results generalise the Bistritz polynomials given in Sec. VI for real parameters (and complex eigenvalues $\lambda$). We refer to that section for more details. More details on the Bistritz algorithm can be found in Bi2002 .
For arbitrary complex parameters and eigenvalue $\lambda$, the polynomial $T_{4}(x)$ reads